OpenStudy (anonymous):

Can someone tell me what program or code you use to insert equations into answers please?

OpenStudy (anonymous):

You can use the equation editor, or you can just type in the latex yourself.

OpenStudy (anonymous):

For example: *[ 5x^3 = \frac{1}{4}(x+9) *] Replace the * with \

OpenStudy (anonymous):

Use the equation editor and type frac{3}{4} I will produce \[\frac{3}{4} \]

OpenStudy (amistre64):

\[\color{#ff0032}{\text{Yep, either way its a pain :)}}\]

OpenStudy (anonymous):

Thanks, lol, I didn't even notice the Equation button, but it is based on LaTeX which I am familiar with.

OpenStudy (anonymous):

If you are familiar with latex you can type it directly rather than using the (clunky imo) editor.